# Enhancing Automatic Chord Recognition via Pseudo-Labeling and Knowledge Distillation

### v1

@@ Abstract
Automatic Chord Recognition (ACR) is constrained by the scarcity of aligned chord labels, as well-aligned annotations are costly to acquire. At the same time, open-weight pre-trained models are currently more accessible than their proprietary training data. In this work, we present a two-stage training pipeline that leverages pre-trained models together with unlabeled audio. The proposed method decouples training into two stages. In the first stage, we use a pre-trained BTC model as a teacher to generate pseudo-labels for over 1,000 hours of diverse unlabeled audio and train a student model solely on these pseudo-labels. In the second stage, the student is continually trained on ground-truth labels as they become available, with selective knowledge distillation (KD) from the teacher applied as a regularizer to prevent catastrophic forgetting of the representations learned in the first stage. In our experiments, two models (BTC, 2E1D) were used as students. In stage 1, using only pseudo-labels, the BTC student achieves over 98% of the teacher’s performance, while the 2E1D model achieves about 96% across seven standard mir_eval metrics. After a single training run for both students in stage 2, the resulting BTC student model surpasses the traditional supervised learning baseline by 2.5% and the original pre-trained teacher model by 1.55% on average across all metrics. And the resulting 2E1D student model improves from the traditional supervised learning baseline by 3.79% on average and achieves almost the same performance as the teacher. Both cases show the large gains on rare chord qualities.
@@ 1 Introduction
Hence, we use unlabeled audio to improve the training of chord recognition and show that the accuracy gains are disproportionately concentrated on rare chord qualities. For this purpose, we propose a two-stage training pipeline (see Figure 3). In the first stage, a pre-trained teacher model generates pseudo-labels for over 1,000 hours of diverse unlabeled audio, and a student model is trained solely on these pseudo-labels until convergence - without requiring any ground-truth annotations. In the second stage, when labeled data becomes available, the pseudo-label-trained student is continually trained on ground-truth labels. To prevent catastrophic forgetting [7] of the representations acquired in the first stage, we apply selective knowledge distillation (KD) from the teacher as a regularizer throughout the second stage.
Prior semi-supervised ACR methods [1, 16] require ground-truth labels from the outset and fuse pseudo-labeled and labeled data together in a single training. Our two-stage training approach decouples labeled and unlabeled data, making the training accessible even when labeled data is initially unavailable. Notably, open-weight pre-trained models are often more readily available than their training data, offering a practical way to leverage teacher knowledge without access to proprietary datasets.
We also introduce a compact, purely transformer-based dual-encoder architecture (2E1D) that is based on [32] and is more lightweight than the teacher model. We demonstrate that our method can generalize across architectures. Our experiments show that the best resulting student surpasses both the traditional supervised learning training method and the pre-trained teacher model across all seven standard mir_eval metrics [26], with particularly large gains on rare chord qualities.
@@ 2 Related Work
Pseudo-labeling, where a trained model generates labels for unlabeled data, has become a cornerstone of semi-supervised learning [15]. The Noisy Student framework [33] demonstrated that iteratively training larger student models on pseudo-labeled data with noise injection can surpass teacher performance, establishing a paradigm for leveraging unlabeled data at scale. FixMatch [28] combined consistency regularization with pseudo-labeling using confidence thresholds, while Mean Teacher [30] introduced exponential moving averages of model weights for stable pseudo-label generation. Meta Pseudo Labels [25] further improved teacher–student training by jointly optimizing the teacher based on student feedback. Within MIR, pseudo-labeling has been successfully applied to various tasks including piano transcription [29], music tagging [14], and speech recognition [18]. For ACR specifically, Bortolozzo et al. [1] adapted Noisy Student for rare chord recognition, employing confidence filtering and iterative teacher–student training to address class imbalance. Li et al. [16] applied contrastive learning to learn chord representations that transfer across datasets. However, these approaches typically require ground-truth labels from the outset and train on mixed pseudo-label and ground-truth signals within a single pipeline. In contrast, we investigate whether separate incremental learning stages can achieve comparable or superior performance by first training on pseudo-labels alone, then adapting to labeled data.
@@
Transferring knowledge via temperature-scaled soft targets, or knowledge distillation (KD) [12], has been shown to provide richer supervision than hard labels. Beyond its original model compression goal, recent work has revealed KD’s regularization properties: soft labels reduce variance at the cost of introducing teacher bias [11], and Yuan et al. [34] studied conditions under which biased soft labels can still improve student generalization. Chen et al. [3] and Mansourian et al. [19] further demonstrated that KD acts as a regularizer against overfitting to noisy annotations. In continual learning, KD has been used to preserve prior knowledge while adapting to new data, mitigating catastrophic forgetting [17, 7]. Learning without Forgetting (LwF) [17] pioneered using distillation from the model’s own previous state to retain old knowledge when learning new tasks. Within MIR, KD has primarily served model compression, training smaller students to match larger teachers for efficient deployment [14]. We apply KD differently: as a regularization mechanism during continual learning that anchors the student to the teacher’s generalized representations. This preserves pseudo-label knowledge when ground-truth labels conflict with teacher predictions, while permitting adaptation when they align. Consequently, KD enables the student to refine decision boundaries without catastrophic forgetting.
@@
Continual learning addresses the challenge of learning from non-stationary data streams without forgetting previously acquired knowledge [31]. In the MIR domain, this challenge is particularly relevant given the ongoing creation of new music and evolving annotation standards. While task-incremental and class-incremental scenarios have received significant attention, data-incremental continual learning remains underexplored for ACR. In this setting, the task and label space remain unchanged but new data arrives.
@@ 3.1 Problem Formulation
Let $\mathcal{D}_{l}=\{(x_{i},y_{i})\}_{i=1}^{N_{l}}$ denote a small labeled dataset where $x_{i}\in\mathbb{R}^{T\times F}$ represents time–frequency features (e.g., Constant-Q Transform) with $T$ frames and $F$ frequency bins, and $y_{i}\in\{1,2,\ldots,C\}^{T}$ are frame-wise chord labels over a vocabulary of size $C$. Let $\mathcal{D}_{u}=\{x_{j}\}_{j=1}^{N_{u}}$ represent large-scale unlabeled datasets where $N_{u}\gg N_{l}$. Our objective is to train an effective student model $f_{s}$ by leveraging pseudo-labels generated from $\mathcal{D}_{u}$. Consequently, this mitigates reliance on expensive manual annotations while maintaining competitive performance.
@@ 3.2 Training Method
We employ a pre-trained teacher model $f_{t}:\mathbb{R}^{T\times F}\rightarrow\mathbb{R}^{T\times C}$ trained on $\mathcal{D}_{l}$ to generate pseudo-labels for unlabeled data. Any open-weight ACR model can serve as the teacher. In our experiments, we use the BTC model of Park et al. [23] as the teacher model, and use the standard vocabulary size $C=170$ for all training settings. For each unlabeled sequence $x_{u}\in\mathcal{D}_{u}$, pseudo-labels are generated via frame-wise argmax over teacher outputs:
$\hat{y}_{u}=\arg\max_{c}\left[f_{t}(x_{u})\right]_{c},\qquad\forall\,\text{frames in }x_{u}$ (1)
@@
Figure 2: Constant-Q Transform (CQT) comparison revealing pitch-shifting artifacts. Top row: original and pitch-shifted (5-5 semitones) spectrograms. Bottom row: artifact intensity maps for $\pm 5$ semitones, computed by realigning shifted CQT bins to compensate for the intended frequency shift.
@@
Figure 3: Illustration of the proposed two-stage training pipeline. Details of audio data are mentioned in 4.1. Note: The resulting Student Model CL of stage 2 can be continually trained when additional labeled data is available.
@@ 3.2.2 The Pipeline of Training
Figure 3 illustrates the complete two-stage training pipeline. In Stage 1, unlabeled data is first preprocessed into Constant-Q Transform spectrograms following the procedure described in Section 4.1. A pre-trained model is used as a teacher to infer frame-wise pseudo-labels (PL) on the unlabeled audio for each track, producing paired data $(Spectrogram,PL)$. A student model is then trained on these pseudo-labeled pairs until convergence. Optionally, knowledge distillation (KD) from the teacher’s soft targets can be applied during this stage to accelerate convergence; this optional path is depicted as a dashed line in the figure. We denote the resulting model after the first stage of training as the Student model PL. Stage 2 begins when the labeled data becomes available. Labeled audio undergoes the same preprocessing, yielding paired data $(spectrogram,GT)$, where GT denotes ground-truth chord annotations. In this stage, Student Model PL serves as the weight initialization for a new model called Student Model CL. Then the Student Model CL serves as the initialization for continual learning as additional labeled datasets are acquired. The selective KD information, mentioned in Section 3.2.4, from the original pre-trained teacher is incorporated throughout the training on Student Model CL. KD acts as regularization, anchoring the student to the teacher’s distributional knowledge when ground-truth labels conflict with teacher predictions, while still allowing adaptation when they agree. In both stages, the training loss is a weighted sum of a classification term and a KD term as follows:
@@ 3.2.4 Selective KD
To further improve robustness, we apply selective KD that filters the teacher signal based on prediction confidence. Let $c=\max_{i}p_{i}^{(t)}$ denote the teacher’s maximum softmax probability. We define an asymmetric weighting function, $w(c)\in[0,1]$ and:
@@
where $\theta_{min}$ filters out uninformative low-confidence samples, $\theta_{max}$ down-weights overconfident predictions that may bias toward majority classes, and factor $K\leq 1$ controls the overconfident predictions rate. The weighted KD loss becomes $\mathcal{L}_{KD}^{sel}=\frac{1}{N}\sum_{i}w(c_{i})\cdot\ell_{KD}(x_{i})$. Samples near decision boundaries (moderate confidence) contain valuable uncertainty information, so we preserve full weight in the informative range $\theta_{min},\theta_{max}]$. We apply selective KD with $\theta_{min}=0.1,\theta_{max}=0.9,K=0.8$ throughout all Stage 2 continual learning experiments to stabilize training by reducing gradient variance from extreme-confidence samples.
@@ 3.2.5 Experimental Models
BTC [23] serves as both the pseudo-labeling teacher and the baseline for self-distillation experiments. It uses a CNN front-end for local spectral feature extraction followed by a bi-directional transformer for temporal modeling, representing a deeper architecture with stacked convolutional and attention layers.
For cross-architecture validation, we introduce a new compact dual-encoder architecture (2E1D) for chord recognition that is purely transformer-based [32] (without any CNN layer). As shown in Figure 4, 2E1D adopts a wider design: (1) a frequency encoder that groups CQT bins into spectral clusters and applies self-attention to learn harmonic relationships across frequency bands; (2) a temporal encoder that processes full-band features to model chord progression patterns over time; and (3) a cross-attention fusion block that combines the two streams for final chord classification. At inference time, we apply a temporal smoothing pipeline over output logits to reduce frame-level prediction jitter. Each class logit channel is convolved with a normalized 1D Gaussian kernel $g[n]=\exp(-n^{2}/2\sigma^{2})/Z$, where $\sigma=k/6$ follows the three-sigma rule for a kernel of width $k$, with replicate padding at segment boundaries. We further process the input using overlapping sliding windows with a stride of $s=\lfloor T(1-r)\rfloor$, where $T$ is the segment length and $r$ is the overlap ratio, accumulating per-class votes across all windows before taking the frame-wise argmax. The use of a wider architecture is to trade depth for parallel processing capacity: the 2E1D model generally processes faster than BTC. We investigate whether attention alone can capture both local spectral patterns and global temporal dependencies without CNN inductive biases [24, 6]. As shown in Section 5, the wider 2E1D architecture is more susceptible to degradation when adapting to noisy labels compared to the deeper architecture of BTC. Thus, the 2E1D model requires stronger KD regularization to maintain stability.  [^1] Models: https://github.com/ptnghia-j/ChordMini
@@ 4.1 Datasets and Preprocessing
Unlabeled Datasets. We use three large-scale unlabeled datasets for pseudo-label generation, totaling over 1,000 hours of audio. The Free Music Archive [4] ($\mathcal{D}_{FMA}$) provides over 100,000 short-form tracks ($\sim$30s) with extensive genre diversity. The MAESTRO dataset [10] ($\mathcal{D}_{MAESTRO}$) contributes 200+ hours of high-quality piano recordings with precise MIDI alignment. The DALI dataset [22] ($\mathcal{D}_{DALI}$) supplies over 5,000 full-length music tracks.
Labeled Datasets. We aggregate annotations from the Isophonics dataset [20], the McGill Billboard corpus [2], RWC Pop [8], and USPop datasets [21], collecting a fixed subset of 600 songs. The dataset is then split with a ratio of 7:1:2 into train/validation/test sets (420/60/120 songs). The “50%” and “full” conditions in our continual learning experiments refer to using 210 and 420 training songs from this subset, respectively.
Clean vs. Noisy Annotations. To evaluate KD’s robustness to annotation noise, we prepare two versions of the labeled data: (1) clean labels with careful manual alignment, and (2) noisy labels sourced online without alignment correction, primarily affecting non-chord (“N”) label boundaries. This setup enables an ablation study of KD’s regularization effect under different noise conditions.
Preprocessing. All audio undergoes identical preprocessing. The Constant-Q Transform (CQT) features [27] are extracted with $F=144$ frequency bins, 24 bins per octave, and a hop length of $h=2048$ samples, yielding a temporal resolution of $\Delta t\approx 93$ms at $f_{s}$= 22.05 kHz. Features undergo z-score normalization using teacher model statistics ($\mu_{t}$, $\sigma_{t}$) to maintain identical input distributions.

### v2

@@ Abstract
Automatic Chord Recognition (ACR) is constrained by the scarcity of aligned chord labels, as well-aligned annotations are costly to acquire. At the same time, open-weight pre-trained models are currently more accessible than their proprietary training data. In this work, we present a two-stage training pipeline that leverages pre-trained models together with unlabeled audio. The proposed method decouples training into two stages. In the first stage, we use a pre-trained BTC model [23] as a teacher to generate pseudo-labels for over 1,000 hours of diverse unlabeled audio and train a student model solely on these pseudo-labels. In the second stage, the student is continually trained on ground-truth labels as they become available. To prevent catastrophic forgetting of the representations learned in the first stage, we apply selective knowledge distillation (KD) from the teacher as a regularizer. In our experiments, two models (BTC, 2E1D) were used as students. In stage 1, using only pseudo-labels, the BTC student achieves over 98% of the teacher’s performance, while the 2E1D model achieves about 96% across seven standard mir_eval metrics. After a single training run for both students in stage 2, the resulting BTC student model surpasses the traditional supervised learning baseline by 2.5% and the original pre-trained teacher model by 1.55% on average across all metrics. The resulting 2E1D student model improves over the traditional supervised learning baseline by 2.67% on average and achieves almost the same performance as the teacher. Both cases show large gains on rare chord qualities.
@@ 1 Introduction
Hence, we use unlabeled audio to improve the training of chord recognition and show that the accuracy gains are disproportionately concentrated on rare chord qualities. For this purpose, we propose a two-stage training pipeline (see Figure 3). In the first stage, a pre-trained teacher model generates pseudo-labels for over 1,000 hours of diverse unlabeled audio, and a student model is trained solely on these pseudo-labels until convergence, without requiring any ground-truth annotations. In the second stage, when labeled data becomes available, the pseudo-label-trained student is continually trained on ground-truth labels. To prevent catastrophic forgetting [7] of the representations acquired in the first stage, we apply selective knowledge distillation (KD) from the teacher as a regularizer throughout the second stage.
Prior semi-supervised ACR methods [1, 16] require ground-truth labels from the outset and fuse pseudo-labeled and labeled data together in a single training run. Our two-stage training approach decouples labeled and unlabeled data, making the training viable even when labeled data is initially unavailable. Notably, open-weight pre-trained models are often more readily available than their training data, offering a practical way to leverage teacher knowledge without access to proprietary datasets.
We also introduce a compact, dual-encoder architecture (2E1D) that is based on the Transformer architecture [32] and is more lightweight than the teacher model. We demonstrate that our method can generalize across architectures. Our experiments show that the best resulting student surpasses both the traditional supervised learning training method and the pre-trained teacher model across all seven standard mir_eval metrics [26], with particularly large gains on rare chord qualities. A chord recognition web application is open-sourced  [^1] https://github.com/ptnghia-j/ChordMiniApp, available for practitioners to experiment with and validate the performance of our models with their music audio locally.
@@ 2 Related Work
Pseudo-labeling, where a trained model generates labels for unlabeled data, has become a cornerstone of semi-supervised learning [15]. The Noisy Student framework [33] demonstrated that iteratively training larger student models on pseudo-labeled data with noise injection can surpass teacher performance, establishing a paradigm for leveraging unlabeled data at scale. FixMatch [28] combined consistency regularization with pseudo-labeling using confidence thresholds, while Mean Teacher [30] introduced exponential moving averages of model weights for stable pseudo-label generation. Meta Pseudo Labels [25] further improved teacher–student training by jointly optimizing the teacher based on student feedback. Within MIR, pseudo-labeling has been successfully applied to various tasks including piano transcription [29], music tagging [14], and more broadly, speech recognition [18]. For ACR specifically, Bortolozzo et al. [1] adapted Noisy Student for rare chord recognition, employing confidence filtering and iterative teacher–student training to address class imbalance. Li et al. [16] applied contrastive learning to learn chord representations that transfer across datasets. However, these approaches typically require ground-truth labels from the outset and train on mixed pseudo-label and ground-truth signals within a single pipeline. In contrast, we investigate whether separate incremental learning stages can achieve comparable or superior performance by first training on pseudo-labels alone, then adapting to labeled data.
@@
Transferring knowledge via temperature-scaled soft targets, or knowledge distillation (KD) [12], has been shown to provide richer supervision than hard labels. Beyond its original model compression goal, recent work has revealed KD’s regularization properties: soft labels reduce variance at the cost of introducing teacher bias [11]. Yuan et al. [34] studied conditions under which biased soft labels can still improve student generalization. Chen et al. [3] and Mansourian et al. [19] further demonstrated that KD acts as a regularizer against overfitting to noisy annotations. In continual learning, KD has been used to preserve prior knowledge while adapting to new data, mitigating catastrophic forgetting [17, 7]. Learning without Forgetting (LwF) [17] pioneered using distillation from the model’s own previous state to retain old knowledge when learning new tasks. Within MIR, KD has primarily served model compression, training smaller students to match larger teachers for efficient deployment [14]. We apply KD differently: as a regularization mechanism during continual learning that anchors the student to the teacher’s generalized representations. This preserves pseudo-label knowledge when ground-truth labels conflict with teacher predictions, while permitting adaptation when they align. Consequently, KD enables the student to refine decision boundaries without catastrophic forgetting.
@@
Continual learning addresses the challenge of learning from non-stationary data streams without forgetting previously acquired knowledge [31]. In the MIR domain, this challenge is particularly relevant given the ongoing creation of new music and evolving annotation standards. While task-incremental and class-incremental scenarios have received significant attention, data-incremental continual learning remains underexplored for ACR. In this setting, the task and label space remain unchanged but new data arrive.
@@ 3.1 Problem Formulation
Let $\mathcal{D}_{l}=\{(x_{i},y_{i})\}_{i=1}^{N_{l}}$ denote a small labeled dataset where $x_{i}\in\mathbb{R}^{T\times F}$ represents time–frequency features (e.g., Constant-Q Transform) with $T$ frames and $F$ frequency bins, and $y_{i}\in\{1,2,\ldots,C\}^{T}$ are frame-wise chord labels over a vocabulary of size $C$. Let $\mathcal{D}_{u}=\{x_{j}\}_{j=1}^{N_{u}}$ represent large-scale unlabeled datasets where $N_{u}\gg N_{l}$. Our objective is to train an effective student model $f_{s}$ by leveraging pseudo-labels generated from $\mathcal{D}_{u}$. This mitigates reliance on expensive manual annotations while maintaining competitive performance.
@@ 3.2 Training Method
We employ a pre-trained teacher model $f_{t}:\mathbb{R}^{T\times F}\rightarrow\mathbb{R}^{T\times C}$ trained on $\mathcal{D}_{l}$ to generate pseudo-labels for unlabeled data. Any open-weight ACR model can serve as the teacher. In our experiments, we use the BTC model of Park et al. [23] as the teacher model, and adopt the standard vocabulary size $C=170$ for all training settings. For each unlabeled sequence $x_{u}\in\mathcal{D}_{u}$, pseudo-labels are generated via frame-wise argmax over teacher outputs:
$\hat{y}_{u}^{(t)}=\arg\max_{c\in\{1,\ldots,C\}}\bigl[f_{t}(x_{u})\bigr]_{t,c},\qquad t=1,\ldots,T$ (1)
@@
Figure 2: Constant-Q Transform (CQT) comparison revealing pitch-shifting artifacts. Top row: original and pitch-shifted (5-5 semitones) spectrograms. Bottom row: artifact intensity maps for $\pm 5$ semitones, computed by realigning shifted CQT bins to compensate for the intended frequency shift. Annotations indicate spectral spreading (low-frequency energy diffusion) and spectral artifacts (high-frequency noise) introduced by the phase-vocoder shift.
@@
Figure 3: Illustration of the proposed two-stage training pipeline. Details of audio data are described in Section 4.1. Note: The resulting Student model CL of stage 2 can be continually trained when additional labeled data is available.
@@ 3.2.2 The Pipeline of Training
Figure 3 illustrates the complete two-stage training pipeline. In Stage 1, unlabeled data is first preprocessed into Constant-Q Transform spectrograms following the procedure described in Section 4.1. A pre-trained model is used as a teacher to infer frame-wise pseudo-labels (PL) on the unlabeled audio for each track, producing paired data $(\text{Spectrogram},\text{PL})$. A student model is then trained on these pseudo-labeled pairs until convergence. Optionally, knowledge distillation (KD) from the teacher’s soft targets can be applied during this stage to accelerate convergence; this optional path is depicted as a dashed line in the figure. We denote the resulting model after the first stage of training as the Student model PL. Stage 2 begins when the labeled data becomes available. Labeled audio undergoes the same preprocessing, yielding paired data $(\text{spectrogram},\text{GT})$, where GT denotes ground-truth chord annotations. In this stage, Student model PL serves as the weight initialization for a new model called Student model CL. Then the Student model CL serves as the initialization for continual learning as additional labeled datasets are acquired. The selective KD signal, mentioned in Section 3.2.4, from the original pre-trained teacher is incorporated throughout the training on Student model CL. KD acts as regularization, anchoring the student to the teacher’s distributional knowledge when ground-truth labels conflict with teacher predictions, while still allowing adaptation when they agree. In both stages, the training loss is a weighted sum of a classification term and a KD term as follows:
@@ 3.2.4 Selective KD
To further improve robustness, we introduce selective KD that filters the teacher signal based on prediction confidence. Let $c=\max_{i}p_{i}^{(t)}$ denote the teacher’s maximum softmax probability. We define an asymmetric weighting function, $w(c)\in[0,1]$ as:
@@
where $\theta_{min}$ filters out uninformative low-confidence samples, $\theta_{max}$ down-weights overconfident predictions that may bias toward majority classes, and factor $K\leq 1$ controls how overconfident predictions are down-weighted. The weighted KD loss becomes $\mathcal{L}_{KD}^{sel}=\frac{1}{N}\sum_{i}w(c_{i})\cdot\ell_{KD}(x_{i})$. Samples near decision boundaries (moderate confidence) contain valuable uncertainty information, so we preserve full weight in the informative range $\theta_{min},\theta_{max}]$. We apply selective KD throughout all Stage 2 continual learning experiments to stabilize training by reducing gradient variance from extreme-confidence samples. We set the hyperparameters with $\theta_{min}=0.1,\theta_{max}=0.9,K=0.8$ as robust defaults based on the observed distribution of the teacher confidence.
@@ 3.2.5 Experimental Models
BTC [23] serves as both the pseudo-labeling teacher and the baseline for self-distillation experiments. The model uses a direct frame projection followed by a stack of bi-directional transformer layers with position-wise feed-forward blocks, representing a deeper architecture with sequentially stacked attention layers.
For cross-architecture validation, we introduce a new compact dual-encoder architecture (2E1D)  [^2] Model: https://github.com/ptnghia-j/ChordMini for chord recognition that is purely transformer-based [32] (without any CNN layer). As shown in Figure 4, 2E1D adopts a wider design: (1) a frequency encoder that groups CQT bins into spectral clusters and applies self-attention to learn harmonic relationships across frequency bands; (2) a temporal encoder that processes full-band features to model chord progression patterns over time; and (3) a cross-attention fusion block that combines the two streams for final chord classification. At inference time, we apply a temporal smoothing pipeline over output logits to reduce frame-level prediction jitter. Each class logit channel is convolved with a normalized 1D Gaussian kernel $g[n]=\exp(-n^{2}/2\sigma^{2})/Z$, where $\sigma=k/6$ follows the three-sigma rule for a kernel of width $k$, with replicate padding at segment boundaries. We further process the input using overlapping sliding windows with a stride of $s=\lfloor T(1-r)\rfloor$, where $T$ is the segment length and $r$ is the overlap ratio, accumulating per-class votes across all windows before taking the frame-wise argmax.
The wider architecture design trades depth for parallel processing capacity: the 2E1D model generally runs faster than BTC. We investigate whether attention alone can capture both local spectral patterns and global temporal dependencies without CNN inductive biases [24, 6]. As shown in Section 5, the wider 2E1D architecture is more susceptible to degradation when adapting to noisy labels compared to the deeper architecture of BTC. Thus, the 2E1D model requires stronger KD regularization to maintain stability.
@@ 4.1 Datasets and Preprocessing
Unlabeled Datasets. We use three large-scale unlabeled datasets for pseudo-label generation, totaling over 1,000 hours of audio. The Free Music Archive [4] ($\mathcal{D}_{FMA}$) provides over 100,000 short-form tracks ($\sim$30s) with extensive genre diversity. The MAESTRO dataset [10] ($\mathcal{D}_{MAESTRO}$) contributes over 200 hours of high-quality piano recordings with precise MIDI alignment. The DALI dataset [22] ($\mathcal{D}_{DALI}$) supplies over 5,000 full-length music tracks.
Labeled Datasets. We aggregate annotations from the Isophonics dataset [20], the McGill Billboard corpus [2], the RWC Pop [8], and the USPop datasets [21], collecting a fixed subset of 600 songs. The dataset is then split in a ratio of 7:1:2 into train/validation/test sets (420/60/120 songs). The “50%” and “full” conditions in our continual learning experiments refer to using 210 and 420 training songs from this subset, respectively.
Clean vs. Noisy Annotations. To evaluate KD’s robustness to annotation noise, we prepare two versions of the labeled data: (1) clean labels with careful manual alignment, and (2) noisy labels sourced online without alignment correction, which primarily affect non-chord (“N”) label boundaries. This setup enables an ablation study of KD’s regularization effect under different noise conditions.
Preprocessing. All audio undergoes identical preprocessing. The Constant-Q Transform features [27] are extracted with $F=144$ frequency bins, 24 bins per octave, and a hop length of $h=2048$ samples, yielding a temporal resolution of $\Delta t\approx 93$ms at $f_{s}=$22.05\text{\,}\mathrm{kHz}$\$. Features undergo z-score normalization using teacher model statistics ($\mu_{t}$, $\sigma_{t}$) to maintain identical input distributions.